# GSCE: A Prompt Framework with Enhanced Reasoning for Reliable LLM-driven Drone Control

### v1

@@ II-B Prompt Engineering
Recent research has increasingly focused on leveraging prompt engineering [18] strategies to enhance the reasoning capabilities of LLMs for robotics. In ChatGPT for Robotics [9], LLMs are provided with a description of guidelines, constraints, and accessible APIs, then LLMs generate code to control robots completing the given task description. Similarly, ViLaIn [19] provides detailed descriptions of the robot and its environmental states, helping LLMs develop a general understanding of the robot’s condition before generating control code. However, LLMs exhibit sensitivity to input perturbations [20, 21, 22]. To address this, PluginSafety [23] applies safety chips that infer specification constraints, ensuring that LLM-generated content adheres to predefined NL constraints. AutoTamp [24] adopts a different approach by prompting the LLM to first generate an intermediate task representation and then translate it into task plans. Additionally, some studies integrate human or automated checker feedback [24, 25, 26, 27] to re-prompt the LLM, refining its outputs for improved reasoning accuracy.
@@ IV Design of GSCE
Our key insight is to develop a robust method that enhances reasoning capabilities for LLM in drone operations. In contrast to previous work that only prompted limited context to LLM [10, 9, 16], our approach explicitly incorporates the usage of constraints within the examples to enhance LLM reasoning capabilities. This method enables the LLM to develop a comprehensive understanding of the constraints and their application, thereby enhancing their ability to reason and apply these constraints in new tasks appropriately. To achieve this, we introduce a structured prompt framework, referred to as GSCE, which consists of guidelines, robotic skill APIs, constraints, and examples. When given an NL task description, the LLM is first prompted with the GSCE framework before generating executable code. The following subsections provide a detailed explanation of the GSCE Prompt Framework.
@@ IV-A GSCE Prompt Framework Overview
  • Robot skills APIs provide detailed instructions on the available APIs that the LLM can utilize when generating code. These instructions include names of APIs and explanations of their functionalities, allowing the LLM to employ the appropriate functions for a given task.
@@ IV-E2 CoT Reasoning
CoT reasoning is a popular prompting strategy that enables LLM to perform reasoning tasks in a step-by-step approach. Given the nature of robotic tasks that are sequentially executed actions, integrating CoT prompting enhances LLM reasoning ability by decomposing a robotic task into a chain of steps. In the GSCE examples, the task is decomposed into sequential movement steps, with the code solutions structured to align with each step. Additionally, the reasoning progress of each step of movement is documented as comments elucidating how each step of thought is grounded. This approach enables LLM to reason step-by step as demonstrated in examples and apply CoT reasoning when given a new task. As shown in Fig.2, LLM imitates illustrations from examples (green block), which effectively decompose new tasks into discrete steps that correspond to a sequence of actions $l=(a_{1}=\text{fly down 5 meters},a_{2}=\text{fly up 4 meters})$.
@@ V-A1 Experimental Environment
The experiment is conducted on a quadcopter using “simple_flight” flight controller [35] in a simulated environment AirSim [36]. AirSim provides realistic environments for testing and developing autonomous systems like drones and self-driving cars, making it well-suited for this research. All tasks were experimented with in the pre-built “Block” environment [37], which features a grid-like layout with basic geometric structures. During the experiment, drone state information was accessible from the simulator and utilized for evaluation purposes.
@@ V-A2 Tasks
Unlike prior studies that primarily focused on simple drone movements (e.g. fly to a position), the proposed tasks comprehensively evaluate the reasoning capabilities of the LLM by encompassing a wide level of complexities. Each task necessitates step-by-step reasoning to analyze the spatial relationships between the target and initial drone states.
@@ V-B1 Metrics
The performance of the abovementioned methods is evaluated using two key metrics: Success Rate (SR) and Completeness. Unlike prior studies [33, 23], which primarily assess whether the task’s final goal is reached, our Success Rate considers all goal states, including intermediate states and the final state. In the real-world scenario, the drone should fly the trajectory precisely as human instructed. Therefore, a task is counted as successful only when the drone reaches the final state while following the correct sequence of intermediate state transitions. The Completeness quantifies the percentage of goal states successfully reached, calculated as the ratio of correct state transitions to the total number of required state transitions in each task. Completeness examines the percentage of intermediate state transitions that the LLM reasoned correctly, offering insights into its reasoning capability across all goal states.

### v2

@@ II-B Prompt Engineering
Recent research has increasingly focused on leveraging prompt engineering [18] strategies to enhance the reasoning capabilities of LLMs for robotics. In ChatGPT for Robotics [9], LLMs are provided with a description of guidelines, constraints, and accessible APIs, and then LLMs generate code to control robots completing the given task description. Similarly, ViLaIn [19] provides detailed descriptions of the robot and its environmental states, helping LLMs develop a general understanding of the robot’s condition before generating control code. However, LLMs exhibit sensitivity to input perturbations [20, 21, 22]. To address this, PluginSafety [23] applies safety chips that infer specification constraints, ensuring that LLM-generated content adheres to predefined NL constraints. AutoTamp [24] adopts a different approach by prompting the LLM to first generate an intermediate task representation and then translate it into task plans. Additionally, some studies integrate human or automated checker feedback [24, 25, 26, 27] to re-prompt the LLM, refining its outputs for improved reasoning accuracy.
@@ IV Design of GSCE
Our key insight is to develop a robust method that enhances reasoning capabilities for LLM in drone operations. In contrast to previous work that only prompted limited context to LLM [10, 9, 16], our approach explicitly incorporates the usage of constraints within the examples to enhance LLM reasoning capabilities. This method enables the LLM to develop a comprehensive understanding of the constraints and their application, thereby enhancing the LLM’s ability to reason and apply these constraints in new tasks appropriately. To achieve this, we introduce a structured prompt framework, referred to as GSCE, which consists of guidelines, robotic skill APIs, constraints, and examples. When given an NL task description, the LLM is first prompted with the GSCE framework before generating executable code. The following subsections provide a detailed explanation of the GSCE Prompt Framework.
@@ IV-A GSCE Prompt Framework Overview
  • Robot skills APIs provide detailed instructions on the available APIs that the LLM can utilize when generating code. These instructions include the names of APIs and explanations of their functionalities, allowing the LLM to employ the appropriate functions for a given task.
@@ IV-E2 CoT Reasoning
CoT reasoning is a popular prompting strategy that enables LLM to perform reasoning tasks in a step-by-step approach. Given the nature of robotic tasks that are sequentially executed actions, integrating CoT prompting enhances LLM reasoning ability by decomposing a robotic task into a chain of steps. In the GSCE examples, the task is decomposed into sequential movement steps, with the code solutions structured to align with each step. Additionally, the reasoning progress of each step of movement is documented as comments elucidating how each step of thought is grounded. This approach enables LLM to reason step-by-step as demonstrated in examples and apply CoT reasoning when given a new task. As shown in Fig.2, LLM imitates illustrations from examples (green block), which effectively decompose new tasks into discrete steps that correspond to a sequence of actions $l=(a_{1}=\text{fly down 5 meters},a_{2}=\text{fly up 4 meters})$.
@@ V-A1 Experimental Environment
The experiment is conducted on a quadcopter using “simple_flight” flight controller [35] in a simulated environment AirSim [36]. AirSim provides realistic environments for testing and developing autonomous systems such as drones and self-driving cars, making it well-suited for this research. All tasks were experimented with in the pre-built environment named “Block”[37]. It has a large flying area with several cubic structures, providing users with an accessible platform for basic testing and execution of drone control scripts. During the experiment, drone state information was accessible from the simulator and utilized for evaluation purposes.
@@ V-A2 Tasks
Unlike prior studies that primarily focused on simple drone movements (e.g., fly to a position), the proposed tasks comprehensively evaluate the reasoning capabilities of the LLM by encompassing a wide level of complexities. Each task necessitates step-by-step reasoning to analyze the spatial relationships between the target and initial drone states.
@@ V-B1 Metrics
The performance of the abovementioned methods is evaluated using two key metrics: Success Rate (SR) and Completeness. Unlike prior studies [33, 23], which primarily assess whether the task’s final goal is reached, our Success Rate considers all goal states, including intermediate states and the final state. In the real-world scenario, the drone should fly the trajectory precisely as human instructed. Therefore, a task is counted as successful only when the drone reaches the final state while following the correct sequence of intermediate state transitions of the ground truth trajectory. The Completeness quantifies the percentage of goal states successfully reached, calculated as the ratio of correct state transitions to the total number of required state transitions of the ground truth trajectory in each task. Completeness examines the percentage of intermediate state transitions the LLM reasoned correctly, offering insights into its reasoning capability across all goal states.